# Statistical Learning of Distributionally Robust Stochastic Control in Continuous State Spaces

### v1

@@ 4 Lower Bounds on Finite Sample Minimax Risks
The constants $C_{1}(k,\delta)$ and $C_{2}(k,\delta)$ are given in the proof in Appendix D.3. This matches the convergence rate up to a $\sqrt{\log n}$ in Theorem 2 and 4. Again, estimating the value function at one single $x$ has the same lower bound.
5 Applications
In this section, we map two applications in finance and engineering into our DRSC framework. Specifically, Section 5.1 outlines a multi-period portfolio optimization problem where market returns may react adversarially to specific actions. Section 5.2 examines a service optimization problem in make-to-order systems, where the distributions of inter-arrival times may be misspecified.
@@ 5.1 Portfolio Optimization
We manage a portfolio of $m$ assets where the time is divided into discrete time periods $t=1,2,...$ (not necessarily uniformly spaced in real time). $X_{t}\in\mathbb{R}^{m}$ denotes the portfolio (or vector of positions) at time $t$, where $X_{t,i}$ is the dollar value of asset $i$ at the beginning of time period $t$: $X_{t,i}>0$ and $X_{t,i}<0$ mean a long position and a short position in asset $i$, respectively.
The dollar value is computed using the current reference price for each asset, which can differ from the current bid or ask price. A reasonable choice is the mid-price between bid and ask [^3] Other, more refined choices exist, such as the volume-weighted average of the bid and ask prices.. One of the assets can be the risk-free asset (i.e. cash), and for convenience designate asset $n$ as the riskless asset (i.e., cash), initially assumed to be non-negative.
We can buy and sell assets at the beginning of each time period. Let $A_{t}\in\mathbb{R}^{m},t=0,1,\dots$ be our decision variables at $t$, which is the dollar values of the trades: $A_{t,i}>0$ or $A_{t,i}<0$ means buying or selling asset $i$ at the beginning of time period $t$, respectively. For simplicity, we assume that there is no contribution of capital. However, one can consume the portfolio by having a total sale that is higher than the total purchase; i.e. $\mathbf{1}^{\top}A_{t}\leq 0$ where $\mathbf{1}\in\mathbb{R}^{n}$ is the column vector of all 1’s. Therefore, the consumption is $C_{t}=-\mathbf{1}^{\top}A_{t}\geq 0$, which is the cash amount taken out from the portfolio. As such, the self-financing constraint (in the absence of any transaction costs) becomes $\mathbf{1}^{T}A_{t}+C_{t}=0$. Hence, it suffices to consider the action space $\mathbb{A}:=\left\{{a\in\mathbb{R}^{m}:\mathbf{1}^{\top}a\leq 0}\right\}$, and $\mathcal{Q}=\mathcal{P}(\mathcal{A})$, as the consumption $C_{t}=-\mathbf{1}^{T}A_{t}$ is determined by a feasible investment $A_{t}$.
At the beginning of the next period, we have $X_{t+1}=W_{t}(X_{t}+A_{t})$, where $W_{t}=\text{diag}(R_{t})\in\mathbb{R}^{m\times m}$ is the matrix of asset returns, and $R_{t,i}$ is the return of the $i$-th asset from period $t$ to period $t+1$.
We assume that the decision maker is endowed with a utility function $U:\mathbb{R}_{+}\rightarrow\mathbb{R}$, which is concave and non-decreasing. Then, taking action $A_{t}$ will incur an instantaneous reward utility $r(X_{t},A_{t})=U(-\mathbf{1}^{\top}A_{t})$. The performance of the policy is measured in terms of the expected infinite horizon discounted total utilities.
As motivated before, in a dynamic portfolio optimization environment, there might be a non-trivial (typically adversarial) market response in reaction to the change in portfolio position induced by current action $A_{t}$. Therefore, a CAA adversary could be a reasonable model for robust dynamic decision-making in this context. Thus, given i.i.d. data $\left\{{R_{t}\in\mathbb{R}^{m},t=1,\cdots n}\right\}$, we build can build Wasserstein distance or $f_{k}$-divergence ambiguity set $\widehat{\mathcal{P}}$ from the empirical measure $\hat{\mu}$ induced by the data. Then, we estimate the DRSC control value by solving (2.2) with $\widehat{\mathcal{P}}$.
5.2 Service and Manufacturing Systems
Following Example 2, we consider a make-to-order systems. The goal is to minimize discounted sum of Waiting time by dynamically adjusting the service rate. The cost of service rates is denoted as $c(A_{n})$, where $c(\cdot)$ is an increasing function. Therefore, the reward function can be written as $r(X_{n},A_{n})=-X_{n}-c(A_{n}).$ And recall the state recursion is $X_{n+1}=(X_{n}+1/{A_{n}}-W_{n})^{+}.$
Here the randomness came from the inter-arrival times $\{W_{n}\}$. The standard queueing theory assumes the demand $W_{1},W_{2},\ldots,W_{n},\ldots$ are i.i.d.. In practice, however, the distribution of $W_{n}$ is often unknown, and the demand sequence $W_{1},W_{2},\ldots,W_{n},\ldots$ is non-i.i.d. and non-stationary. For example, an empirical study (Kim and Whitt, 2014) and a publicly available dataset from a US bank call center, ”DataMOCCA” (Trofimov et al., 2006), indicate that arrivals at call centers and hospitals exhibit significant time-of-day and day-of-week effects. Therefore, it is important to model this ambiguity using our distributionally robust framework. In this framework, we use adversaries to model all possible inter-arrival (joint) distributions the system manager may face. It is natural to use a CAU adversary, as this inter-arrival time should not depend on the manager’s service decision.

### v2

@@ 5 Learning Optimal DRSC: Minimax Risks Lower Bounds
The constants $C_{1}(k,\delta)$ and $C_{2}(k,\delta)$ are given in the proof in Appendix D.3. This matches the convergence rate up to a $\sqrt{\log n}$ in Theorem 3 and 5. Again, estimating the value function at one single $x$ has the same lower bound.
Therefore, we have shown that even for the simple DRSC instance in Lemma 1, the minimax risk admits lower bounds that match the convergence rate upper bounds in Theorems 1–4. This establishes the tightness of our results and verifies the rates reported in Table 1.
6 Algorithm Design
Recall from our theoretical analysis that the solution $\hat{u}$ of the empirical Bellman equations achieves minimax-optimal convergence rates in estimating the true robust optimal values $u^{\ast}$ and $\bar{u}$. Motivated by this result, our algorithms are designed to approximate the empirical Bellman equations, $\mathbf{T}(u)=u$ and $\overline{\mathbf{T}}(u)=u$. In what follows, we focus on the $f_{k}$-divergence ambiguity setting and develop actor-critic algorithms with neural-network function approximation for DRSC under both CAA and CAU adversary settings.
A key advantage of our approach is its seamless integration into standard RL pipelines: robustness can be introduced into existing actor-critic frameworks with only modest adjustments, without the need for extensive redevelopment. The method retains the familiar modular structure of separate neural policy and value function approximators, and can be implemented with only minor changes to widely used architectures and training loops. As a result, it is readily compatible with popular RL frameworks and easily adaptable across a wide range of domains.
6.1 The CAA Case
In the CAA setting, we parameterize the value function $u_{\theta}:\mathbb{X}\rightarrow\mathbb{R}$ and the policy $\pi_{\eta}:\mathbb{X}\rightarrow\mathbb{A}$ as neural networks. Leveraging the actor-critic framework, we alternate between Bellman error minimization and policy improvement steps, illustrated by Algorithm 1. Specifically, for a given policy-value pair $(\pi_{\eta},u_{\theta})$, the empirical Bellman operator is defined as:
$\mathbf{T}_{\eta,\theta}(x):=\inf_{\psi\in\widehat{\mathcal{P}}}\int_{\mathbb{W}}r(x,\pi_{\eta}(x),w)+\alpha u_{\theta}(f(x,\pi_{\eta}(x),w))\psi(dw),$
where the minimization is performed over the empirical $f_{k}$-divergence ambiguity set $\widehat{\mathcal{P}}$ centered around the empirical measure $\hat{\mu}$, as described in previous sections.
Strong Duality and Optimal Lagrange Multiplier: To compute gradients of the empirical Bellman operator, we invoke strong duality [9], transforming the minimization over probability measures into an equivalent one-dimensional convex problem:
$\mathbf{T}_{\eta,\theta}(x)=\sup_{\lambda\in\mathbb{R}}\bigg\{\lambda-c_{k}(\delta)\left[\int_{\mathbb{W}}(r(x,\pi_{\eta}(x),w)+\alpha u_{\theta}(f(x,\pi_{\eta}(x),w))-\lambda)_{+}^{k^{\prime}}\hat{\mu}(dw)\right]^{1/k^{\prime}}\bigg\}.$
Here, $\lambda$ is the dual multiplier. Under mild assumptions, the maximization in $\lambda$ has a unique optimal solution $\lambda^{*}(\eta,\theta,x)$, which can be efficiently computed using bisection search. We define the intermediate operator for convenience:
$\widehat{\mathbf{T}}_{\eta,\theta}(\lambda,x)=\lambda-c_{k}(\delta)\left[\int_{\mathbb{W}}(r(x,\pi_{\eta}(x),w)+\alpha u_{\theta}(f(x,\pi_{\eta}(x),w))-\lambda)_{+}^{k^{\prime}}\hat{\mu}(dw)\right]^{1/k^{\prime}}.$
Bellman Error Minimization: Given a fixed policy $\pi_{\eta}$, we minimize the squared $L^{2}$ Bellman error:
$\min_{\theta}\int_{\mathbb{X}}[u_{\theta}(x)-\mathbf{T}_{\eta,\theta}(x)]^{2}\nu(dx),$
where $\nu$ is a user-specified probability measure supported on the entire state space $\mathbb{X}$. To compute gradients, we utilize the envelope theorem to get that:
$\nabla_{\theta}\mathbf{T}_{\eta,\theta}(x)=\nabla_{\theta}\widehat{\mathbf{T}}_{\eta,\theta}(\lambda^{*}(\eta,\theta,x),x).$ (6.1)
To minimize Bellman Error we update the $\theta$ using first-order algorithms. For illustrative purposes, we formulate the mini-batch stochastic gradient descent in this context. At each iteration, we independently sample states $X_{i}\sim\nu$, compute the optimal multipliers $\lambda_{i}^{*}$ in parallel, and update the parameters as follows:
$\theta_{t+1}=\theta_{t}-\ell_{t}\frac{1}{n}\sum_{i=1}^{n}\nabla_{\theta}[(u_{\theta}(X_{i})-\mathbf{T}_{\eta,\theta}(x))^{2}],$ (6.2)
where the gradient w.r.t. $\theta$ is computed using the chain-rule and (6.1).
Policy Improvement: Given a fixed value function $u_{\theta}$, the policy $\pi_{\eta}$ is improved by solving the following optimization problem with first-order methods:
$\max_{\eta}\int_{\mathbb{X}}\mathbf{T}_{\eta,\theta}(x)\nu(dx)$
with gradients computed similarly using the envelope theorem. Again, we update parameters via stochastic gradient ascent:
$\eta_{t+1}=\eta_{t}+\ell^{\prime}_{t}\frac{1}{n}\sum_{i=1}^{n}\nabla_{\eta}\widehat{\mathbf{T}}_{\eta,\theta}(\lambda_{i}^{*},X_{i}).$ (6.3)
Algorithm 1 Robust Actor–Critic Algorithm for CAA
1:Step–size schedules $\{\ell_{t}\}_{t\geq 0},\{\ell^{\prime}_{t}\}_{t\geq 0}$, batch size $n$, initial parameters $(\theta_{0},\eta_{0})$, distribution $\nu$ on $\mathbb{X}$
2:Repeat until convergence:
3:
4:1. Bellman error minimization (critic) update
5:Sample states $\{X_{i}\}_{i=1}^{n}\overset{\text{i.i.d.}}{\sim}\nu$
6:for $i=1,\dots,n$ in parallel do
7:  $\lambda_{i}^{\!*}\;\leftarrow\;\arg\max_{\lambda\in\mathbb{R}}\;\widehat{\mathbf{T}}_{\eta_{t},\theta_{t}}(\lambda,X_{i})$ $\triangleright$ bisection search
8:end for
9:Update $\theta$ as in Equation (6.2)
10:2. Policy improvement (actor) update
11:Resample (or reuse) $\{X_{i}\}_{i=1}^{n}\sim\nu$
12:for $i=1,\dots,n$ in parallel do
13:  Recompute $\lambda_{i}^{\!*}\leftarrow\arg\max_{\lambda}\widehat{\mathbf{T}}_{\eta_{t},\theta_{t+1}}(\lambda,X_{i})$
14:end for
15:Update $\eta$ as in Equation (6.3)
[…]
17:  return final parameters $(\theta_{t},\eta_{t})$
6.2 The CAU Case with Continuous Action
While our theoretical analysis focuses primarily on finite action models in the CAU setting, practical applications can involve continuous action spaces. Hence, we propose an extended algorithm suitable for continuous state-action problems. The CAU case is inherently more complex since deterministic policies are generally suboptimal, necessitating the consideration of randomized policies. To address this, we propose a novel generative policy approach. Specifically, we view a policy as a generative model that produces randomized actions given a state $x\in\mathbb{X}$ by utilizing an external source of randomness.
Concretely, we define the policy as a mapping $\pi_{\eta}:\mathbb{R}^{d}\times\mathbb{X}\rightarrow\mathbb{A}$, where an action is generated according to $\pi_{\eta}(N,x)$ and $N\sim N(0,I)$ is a standard normal random vector independent of the state. For each state $x$, we sample a normal vector $N$ of $d$ dimensions. Due to space constraints, a detailed discussion of the algorithm for the CAU case is provided in Appendix E.
7 Experiments
We provide two sets of experiments using the DRSC framework with CAA and CAU adversaries proposed above. The first one is a study on inventory control (Example 2 in Section 3.3) with real-world data, and the second one is a simulation study on portfolio optimization (Example 3 in Section 3.3). For each setting, we solve the robust value function and policy using the algorithm described in Section 6, and compare it with the non-robust optimal policy.
7.1 Inventory Control
Inventory management is a cornerstone of retail and supply chain operations. Managers must balance the cost of carrying excess stock against the risk of lost sales when demand outstrips supply. This trade-off becomes particularly challenging when customer demand is uncertain and subject to shifts caused by seasonality, promotions, or broader economic disruptions. Traditional models often assume that the demand process is i.i.d. across time, but in practice, this assumption rarely holds. Unexpected fluctuation, such as unaccounted shifts in consumer preferences or disruptions in the broader economy, can cause substantial performance degradation when managers rely solely on non-robust policies trained under the assumption of i.i.d. demand.
In this section, we empirically demonstrate how managers can benefit from our distributionally robust approach to improve decision-making in a data-driven context. By explicitly accounting for non-stationary shifts and potential adversary behavior in the demand process, our framework enables the design of effective and robust ordering policies that hedge against worst-case scenarios.
7.1.1 Setting
We consider Example 2 in Section 3.3, where a decision maker must determine the quantity of a single product to order over discrete time periods $t=0,1,2,\dots$. Let $X_{t}\in\mathbb{R}_{\geq 0}$ denote the inventory level at the beginning of period $t$, and let $A_{t}\in\mathbb{R}_{\geq 0}$ be the quantity ordered at the beginning of that period. We assume that orders placed at time $t$ arrive instantaneously and are available to satisfy demand in the same period.
During each period, a random demand $W_{t}\in\mathbb{R}_{\geq 0}$ is realized, and unmet demand is either backlogged or lost. For simplicity, we assumed the unmet demand is directly lost.
The performance of a policy is measured via the expected infinite horizon discounted total cost. Note that our theorems can be extended to the case where the reward is the function of $X_{t},A_{t},W_{t}$ as shown before.
We use the open Rossmann Store Sales dataset from Kaggle  [^1] https://www.kaggle.com/competitions/rossmann-store-sales/. It contains the demand data from 2013 to 2016. For one store, we use the earliest 80% data as the training set from which we calculate the empirical distribution of $W_{t}$; we test the trained robust policies on the time sequence of the latest 20% data. To confirm the generalizability of our results, we repeat the same experiment procedure on 5 different stores and report all the results. We set $h=0.1,c=1,p=2,\alpha=0.9$. More implementation details can be found in Appendix F.
7.1.2 Results
The results are shown in Table 2. Note that in this case, the same-day demand does not react to $A_{t}$, and it shows a trend of auto-correlation, so CAU is appropriate. For each CAU policy, we report the variance of the return by repeatedly sampling the normal random sequence 50 times. In each case, we could calculate the explicit solution of the non-robust optimal policy [36], and compare the performance of the CAA and CAU policy with it. We find that
  1. 1.
    For properly small and mild $\delta$, both CAA and CAU robust policies outperform the non-robust optimal policy, which also serves as an illustration of the distribution shift in demand over time.
  2. 2.
    For small $\delta$, CAA outperforms CAU in most cases.
  3. 3.
    For mild and large $\delta$, CAU outperforms CAA, and it also has the hightest returns in 4 out of 5 cases.
Note that the non-robust policy is a base stock policy (S-policy). And we find that the CAA robust policies are also S-policies, which corresponds to the explicit solutions in [44]. We also see a positive relation between the size of the uncertainty set and randomization - the CAU optimal policy becomes more randomized as the uncertainty set get larger. However, the randomization is not always better, especially for small $\delta$, because the distribution shift may not be in the adversarial direction. The plots of CAA and CAU policies are in Appendix F.
Table 2: Return of CAA and CAU policies for 5 stores
Return ($\times 10^{3}$) $\delta=0.001$ $\delta=0.01$ $\delta=0.1$ $\delta=0.5$ $\delta=1$
CAA 38.04 38.38 38.47 37.05 36.61
[…]
$\pi_{0}^{*}$ 37.85
Return ($\times 10^{3}$) $\delta=0.001$ $\delta=0.01$ $\delta=0.1$ $\delta=0.5$ $\delta=1$
CAA 81.72 81.83 81.97 76.48 70.65
[…]
$\pi_{0}^{*}$ 81.60
Return ($\times 10^{3}$) $\delta=0.001$ $\delta=0.01$ $\delta=0.1$ $\delta=0.5$ $\delta=1$
CAA 31.75 32.21 33.17 29.67 29.64
[…]
$\pi_{0}^{*}$ 31.16
Return ($\times 10^{3}$) $\delta=0.001$ $\delta=0.01$ $\delta=0.1$ $\delta=0.5$ $\delta=1$
CAA 62.28 63.12 64.47 62.02 54.70
[…]
$\pi_{0}^{*}$ 62.26
Return ($\times 10^{3}$) $\delta=0.001$ $\delta=0.01$ $\delta=0.1$ $\delta=0.5$ $\delta=1$
CAA 61.20 62.16 63.11 60.98 60.20
[…]
$\pi_{0}^{*}$ 59.66
Our findings highlight several takeaways for managers:
  1. 1.
    Value of Robustness under Demand Shifts. When demand patterns shift over time, robust policies consistently outperform the classical non-robust policy. This underscores the importance of preparing for distributional changes rather than optimizing exclusively for historical distributions.
  2. 2.
    Different Forms of Robustness. For relatively small deviations from historical demand, CAA performs best; for larger deviations, CAU dominates. This suggests that managers should calibrate the level of robustness to the degree of environmental volatility they expect.
  3. 3.
    Trade-off Between Randomization and Performance. We observe that CAU policies introduce more variability in ordering decisions as uncertainty grows. While this can be beneficial when facing severe disruptions, excessive randomization may reduce performance when demand is relatively stable. Hence, robustness should be deployed thoughtfully, balancing protection against downside risk with efficiency in normal times.
@@ 7.2 Portfolio Optimization
In addition to the real data case, we want to test our proposed method on a more challenging case, where several constraints are violated. We find our method still applicable to this hard case with some numerical techniques.
7.2.1 Motivation
Portfolio allocation is one of the most fundamental decisions in finance, requiring investors to balance consumption today against investment for tomorrow. A classical result from the Merton model [35] shows how investors allocate between risk-free and risky assets under known return distributions. However, in practice, asset returns are uncertain, non-stationary, and subject to shocks such as recessions, market bubbles, or geopolitical events. Standard optimization methods that rely on historical averages may therefore expose investors to substantial downside risk.
Our framework addresses this challenge by designing policies that remain effective even when the return distribution shifts. Specifically, we test how distributionally robust policies—modeled under our CAA and CAU approaches—alter portfolio allocation strategies in the face of adversarial shocks. This provides insights into how investors should adjust their risk exposure when market conditions deviate from expectations.
7.2.2 Setting
@@
We consider Example 3 in Section 3.3, where we want to manage a portfolio of $m$ assets where the time is divided into discrete time periods $t=0,1,2,$... (not necessarily uniformly spaced in real time). $X_{t}\in\mathbb{R}^{m}$ denotes the portfolio (or vector of positions) at time $t$, where $X_{t,i}$ is the dollar value of asset $i$ at the beginning of time period $t$: $X_{t,i}>0$ and $X_{t,i}<0$ mean a long position and a short position in asset $i$, respectively. We can buy and sell assets at the beginning of each time period. Let $A_{t}\in\mathbb{R}^{m},t=0,1,\dots$ be the resulting percentage of the $m$ assets after trading at $t$: $A_{t,i}>1$ or $0\leq A_{t,i}<1$ means buying or selling asset $i$ at the beginning of time period $t$, respectively. The performance of the policy is measured in terms of the expected infinite horizon discounted total utilities.
We consider the 2-asset case, i.e. $m=2$. This is a common case where one asset is risk-free while the other is risky. We set the risk-free return to be 1.04, and the risky return distribution to be lognormal with expected return around 1.13; i.e. $r_{f}=1.04,~\psi_{0}(r_{risky})\sim log\mathcal{N}(0,0.5)$.
Assume the reward is the utility from consumption as in Merton model [35], $r(X_{t},A_{t},W_{t})=\frac{C_{t}^{\gamma}}{\gamma}$. We set $\gamma=0.5$, and the discount factor $\alpha=0.9$. Also, we impose another assumption that the total consumption should not exceed the total assets; $\mathbf{1}^{\top}X_{t}-\mathbf{1}^{\top}A_{t}\leq\mathbf{1}^{\top}X_{t}$, i.e. $\mathbf{1}^{\top}A_{t}\geq 0$. And we assume that $\mathbf{1}^{\top}X_{0}\geq 0$ in the beginning.
Note that based on our constraints, once $\mathbf{1}^{\top}X_{t}\leq 0$, we must have $0\leq\mathbf{1}^{\top}A_{t}\leq\mathbf{1}^{\top}X_{t}\leq 0$. Thus, instead of waiting for $\mathbf{1}^{\top}X_{t}$ to become positive again (which may never happen), we end the process immediately.
7.2.3 Results
(a) Slice of CAA value function at $x_{1}=0$
(b) Slice of CAU value function at $x_{1}=0$
Figure 1: Robust value functions with various sizes of the uncertainty set.
(a) Slice of CAA policy at $x_{1}=0$
(b) Slice of CAU policy at $x_{1}=0$
Figure 2: Robust policies with various sizes of the uncertainty set.
Figure 3: Value difference between non-robust optimal policy and two robust policies.
We compare three strategies:
  • the classical non-robust policy based on historical distributions,
  • the robust CAA policy, and
  • the robust CAU policy.
Robust policies are trained under varying levels of uncertainty $\delta$, reflecting different degrees of market stress or volatility. The results are shown in Figure 1, 2 and 3. In Figure 1, $\text{MC}_{\psi_{0},\text{caa},0.0005}$” is the Monte Carlo value estimation of the robust policy with $\delta=0.0005$ in the original environment $\psi_{0}$ (in order to verify the correctness of the converged value and policy).
We have the following observations:
  1. 1.
    The robust value becomes smaller as the size of the uncertainty set $\delta$ increases. For the same $\delta$, the value of CAU policy is higher than that of CAA policy. This is because CAU allows randomized policies which is at least no worse than CAA policy.
  2. 2.
    The robust policy becomes more conservative as the size of the uncertainty set $\delta$ increases, in the sense that more risk-free and less risky asset would be held.
  3. 3.
    We set the real $\delta$ to be $10^{-2}$, and we consider the $\chi^{2}$ divergence. This means that the test distribution of the risky asset is $log\mathcal{N}(-\sqrt{\ln(1+2\delta)}/2,0.5)$. With the adversarial distribution shift, the CAU policy is consistently better than the CAA policy. With small and mild $\delta$, both robust policies are better than the non-robust policy.
As mentioned before, because the trader here is small relative to the market and returns are driven by external factors, it is reasonable to use the CAU framework.
Our analysis yields several finance-relevant takeaways:
  1. 1.
    Flight to safety under heightened uncertainty. As the size of the ambiguity set $\delta$ increases, both robust policies reallocate wealth away from the risky asset toward the risk-free asset. This mirrors the “flight to quality” observed in financial markets during periods of elevated volatility, where investors naturally hedge by holding safer assets. Our results formalize this behavior as the optimal response to distributional ambiguity.
  2. 2.
    Adaptive diversification through CAU policies. The CAU framework, which allows randomized allocations across assets, consistently delivers higher utility than CAA when market shocks are severe. This corresponds to a form of adaptive diversification—spreading exposure in a probabilistic manner rather than committing deterministically to one allocation. In practice, this finding suggests that fund managers facing crisis conditions may benefit from dynamic or mixed strategies that preserve flexibility in the face of adversarial return scenarios.
  3. 3.
    Robustness improves risk-adjusted performance. For small to moderate levels of distributional shifts, both CAA and CAU policies outperform the classical non-robust benchmark. In finance terms, this means that incorporating robustness can improve Sharpe-like efficiency: investors attain higher expected utility (adjusted for risk aversion) without incurring additional downside exposure. Robust optimization thus not only insures against tail events but also enhances performance in moderately perturbed environments.
Together, these insights demonstrate that robust portfolio policies provide a disciplined way to operationalize risk management principles widely practiced in finance—hedging in uncertain markets, diversifying under stress, and improving risk-adjusted returns when market conditions deviate modestly from historical patterns.